# Supervising 3D Talking Head Avatars with Analysis-by-Audio-Synthesis

### v2

@@ 4.2 Speech-Driven Facial Animation Diffusion
Figure 3: Diffusion architecture. The upper part of the figure (green) depicts the architecture of the diffusion model and the lower part (yellow) illustrates the application of the audio consistency loss. Gray boxes indicate the input to the system. Trainable components are highlighted in orange and frozen ones in blue.
Architecture. The core of THUNDER consists of a denoiser in the form of a transformer decoder [vaswani2017attention]. Similar to other contemporary 3D animation methods (e.g. Zhao et al. [zhao2024media2face] and Sun et al. [sun2024diffposetalk]), THUNDER processes a noisy sequence of expression parameters $\boldsymbol{\tilde{x}}_{(d)}^{1:T}$, where $d$ represents the diffusion timestep. These parameters serve as the query inputs to the transformer. The keys and values for the transformer are comprised of the input audio features $\mathbf{a}^{1:T}$ and a diffusion step embedding $\mathcal{T}(d)$, concatenated along the temporal dimension. The function $\mathcal{T}$ is defined as a sinusoidal timestep function similar to Ho et al. [ho2020denoising]. We do not incorporate additional animation-controlling conditions (such as style vectors [cudeiro2019capture, fan2022faceformer, xing2023codetalker], emotion vectors [danecek2023emote, Peng2023_EmoTalk] or CLIP embeddings [zhao2024media2face]) as this is not the goal of this work. However, the design can easily be augmented with additional input conditions. All inputs are projected into the transformer’s feature dimension, $f=128$, using dedicated learnable affine layers. Positional encoding within the transformer’s self-attention utilizes the ALiBi mechanism [press2021train]. The cross-attention layers use a diagonal binary mask to facilitate the attention mechanism, and no additional positional encoding. The model predicts the reconstructed expression parameters $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. The transformation can be formally described by the following equation:
@@
Here, $d\in\{1,\dots,D\}$ indicates the diffusion timestep, and $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$ is the denoised prediction. Unlike some methods that predict noise, this model directly predicts the fully denoised space, which is used to compute the 3D mesh sequence: $M(\boldsymbol{\beta}=\mathbf{0},\boldsymbol{\widehat{x}}_{(0)}^{1:T})\rightarrow\widehat{\mathbf{V}}_{(0)}^{1:T}$. This approach allows the model to define losses directly on the final 3D mesh space, crucial for leveraging mesh-to-speech. See Fig. 3 for an overview.
@@
[…]
$\mathcal{L}_{rec}=$ $\ \textup{MSE}(\mathbf{V}^{1:T},\widehat{\mathbf{V}}_{(0)}^{1:T})+\textup{MSE}(\boldsymbol{\psi}^{1:T},\boldsymbol{\widehat{\psi}}_{(0)}^{1:T})+$
$\ \textup{MSE}(\boldsymbol{\theta}_{jaw}^{1:T},{\boldsymbol{\widehat{\theta}}_{jaw}^{1:T}}_{(0)}).$
[…]
$\mathcal{L}_{vel}=$ $\ \textup{MSE}(v(\mathbf{V}^{1:T}),v(\widehat{\mathbf{V}}_{(0)}^{1:T}))+\textup{MSE}(v(\boldsymbol{\psi}^{1:T}),v(\boldsymbol{\widehat{\psi}}_{(0)}^{1:T}))+$
$\ \textup{MSE}(v(\boldsymbol{\theta}_{jaw}^{1:T}),v({\boldsymbol{\widehat{\theta}}_{jaw}^{1:T}}_{(0)})),$
@@
with $w_{mts}=1$. The model is trained using classifier-free guidance [ho2022cfg]. The audio condition is randomly dropped and replaced with a learnable vector with 20% probability.
Inference. At inference time, we initialize the first step by sampling the Gaussian noise: $\boldsymbol{\widehat{x}}_{(D)}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. We then proceed to denoise for $D$ diffusion steps. In each step, we first employ the denoiser $\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})$ to obtain $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. Then, we compute $\boldsymbol{\tilde{x}}_{(d-1)}$ by adding the expected amount of noise back to obtain $\boldsymbol{\tilde{x}}_{(d-1)}$. Since THUNDER is trained with classifier-free guidance, it is possible to trade-off fidelity of the lip-sync for increased diversity by combining $s_{a}\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})+(1-s_{a})\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\O)$ during the denoising process. However, we find our results are diverse enough without this and so we effectively set $s_{a}=1$.
@@ 5.1 Datasets
The selection of datasets for THUNDER is challenging. On the one hand, 4D scan datasets such as VOCASET [cudeiro2019capture] provide good reconstructions but are too small to train generative models. On the other hand, in order to have a high-enough quality geometry from which speech can be predicted, we cannot rely on in-the-wild datasets, as reconstruction artifacts decrease the quality of the GT and hence our mesh-to-speech and talking head models.
Hence, we use a collection of public in-the-lab captured video datasets commonly used for lip-reading, namely TCD-TIMIT [Harte2015tcdtimit], RAVDESS [livingstone_2018_1188976] and GRID [cooke2016grid], totaling 117 subjects, uttering 42140 short sentences, totaling 3535816 frames at 25FPS (over 39 hours). These datasets provide clear imagery without in-the-wild complexity and can be reconstructed by off-the-shelf methods to a satisfactory degree. Following EMOTE [danecek2023emote], we utilize the publicly available INFERNO framework [inferno2023], which contains a SOTA face reconstruction system [MICA:ECCV2022, EMOCA:CVPR:2021, filntisis2022visual]. We split each of the datasets by subjects, 70% for training and 15% each for validation and testing.
@@ 5.2 Mesh-To-Speech
Implementation details. We initialize the conformer and prediction heads with pretrained weights from Choi et al. [choi2023lip2speechunit]. The input MLP is trained from scratch. We train the model with the Adam optimizer [Kingma2015] for 20 epochs and select the checkpoint with the lowest validation loss (reached within 5 epochs). The training batch size is 16 and the input sequence is trimmed to 125 frames.
Quantitative comparison. Since the mesh-to-speech model is the first of its kind, we compare mesh-to-speech to a recent silent-video-to-speech model [choi2023lip2speechunit] which we finetune on the videos from our training dataset for fair comparison. Note that FLAME does not model the teeth or tongue, which makes the mesh-to-speech task arguably harder than video-to-speech. Teeth and tongue, which may be visible on video, are critical for production of certain phonemes, such as labiodentals (/f/, /v/), or alveolar phonemes (/s/, /z/, /t/, /d/, /l/) [Ladefoged2011phonetics]. Consequently, we do not expect mesh-to-speech to match the performance of a video-based method.
Table 1 reports the standard video-to-speech metrics, namely Short-Time Objective Intelligibility [Taal2011stoi], Extended STOI [Jensen2016estoi] to measure the intelligibility of the generated samples, Perceptual Evaluation of Speech Quality (PESQ) [Rix2001pesq] (narrow and wide band) to measure the perceptual quality, and Word Error Rate (WER), which we measure with the SOTA ASR network Whisper[radford2023whisper]. As expected, mesh-to-speech is not as accurate as video-to-speech, but the performance is comparable, suggesting that there is sufficient audio signal in the 3D to be useful for our downstream task. Note that mesh-to-speech is, however, more accurate than Choi et al.’s original (but not the fine-tuned) model. We experiment with the following input spaces for mesh-to-speech - selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$ (mouth2s), all vertices $\mathbf{V}$ (exp2s), and FLAME expression parameters $\boldsymbol{x}$ (exp2s) and report the results in Tab. 1. Remarkably, the model which takes global expression vectors, performs slightly better than the selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$. Despite that, we choose mouth2s to be our final model, as it results in better lip-sync supervision (discussed in Sec. 5.3 and Tab. 2). Please refer to the Sup. Video for audible comparison and Sec. 7.1 in Sup. Mat. for further discussion on performance.
@@
Mesh-to-speech exp2speech $\boldsymbol{x}$ 0.502 0.273 1.257 1.468 0.648
@@
Video-to-speech Choi et al. (finetuned) 0.555 0.348 1.281 1.511 0.437
@@
Table 1: Quantitative comparison of video-to-speech and forms of mesh-to-speech (mouth2speech, face2speech and exp2speech).
@@ 5.3 Speech-Driven Facial Animation
Implementation details. THUNDER is trained using the Adam optimizer for 260 epochs with batch size 48, sequence length of 70 frames, and $D=1000$ diffusion steps with a linear noise schedule. The transformer has 8 layers with 4 heads and feature dim $f=128$.
Baselines. We compare THUNDER to other speech-driven avatar methods that are also trained on pGT and which output 3DMM parameters (as opposed to vertices). We reimplemented and trained several SOTA methods, namely FlameFormer (an adaptation of FaceFormer [fan2022faceformer]), Media2Face [zhao2024media2face] and DiffPoseTalk [sun2024diffposetalk] (see Sup. Mat. for details). Unless stated otherwise, we keep the pretrained Wav2Vec weights frozen. Models that finetune Wav2Vec are denoted with a suffix -T (T for trainable).
@@
Lip Vertex Error. Following [richard2021meshtalk, sun2024diffposetalk, xing2023codetalker] we report LVE, which calculates the maximum L2 error of all lip vertices for each frame and then computes the average over all frames.
Dynamic Time Warping. We also report the DTW proposed by Thambiraja et al. [bala2022imitator]. First, distances between the mid-points of lower and upper lips are calculated for both the predictions and GT. The two resulting time series are then used to compute the DTW distance.
Lip Correlation Coefficients. We argue that the above metrics are not comprehensive enough as LVE only operates over a sparse set of lip vertices and does not model the temporal aspect very well as each frame is handled separately and DTW only models the lower-upper lip distance. Specifically, we compute Pearson and Concordance correlation coefficients for every vertex coordinate in the mouth-region and average the coefficients. PCC and CCC are commonly used metrics for comparing time series.
Face dynamic deviation. An important aspect of facial animation, is the expressiveness of the upper face vertices. Following CodeTalker [xing2023codetalker], we report Face Dynamics Deviation (FDD), which measures the differences between the temporal standard deviation of the pGT and the predicted vertices, averaged over a set of vertices. We report upper-FDD (FDD-U) and lip-FDD (FDD-L).
Sample diversity. One of the most important aspects of stochastic models is that they generate diverse animations for the same input. Remarkably, previous work has not analyzed this aspect [stan2023facediffuser, sun2024diffposetalk, zhao2024media2face]. We compute standard deviations of vertex distances between the pGT and the predictions for the same input at frame $t$. Formally, given a tensor of per-vertex L2 distances $\mathbf{D}\in\mathbb{R}^{N\times S\ \times T\times n_{v}}$, with $N$ being the size of the test dataset and $S=32$ is the number of sampled outputs per input. We compute the standard deviation along the sample dimension $S$. Other dimensions are subsequently averaged. We measure sample diversity for the upper face region (S-Div-U) and the lip region (S-Div-L).
Temporal diversity metric. We also report the temporal diversity, where the standard deviation is computed over the temporal dimension of $\mathbf{D}\in\mathbb{R}^{N\times S\ \times T\times n_{v}}$ (and averaged across the rest). Again, we calculate the diversity for both lip (T-Div-L) and upper face regions (T-Div-U).
@@ 5.3.1 Quantitative evaluation
Tab. 2 reports the computed metrics computed on the test set for THUNDER, baselines and ablations.
Experiment Model LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ T-FDD-U $\downarrow$ T -FDD-L $\downarrow$
THUNDER and mesh-to-speech input space THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
@@
THUNDER and trainable audio enc. THUNDER-T w/o m2s 0.723 0.428 0.623 0.266 0.02 0.0411 0.0656 0.216 0.0118 0.0811
@@
Mesh-to-speech with other methods and additional input conditions FlameFormer* 0.809 0.368 0.57 0.291 0.0271 0.0372 0.132 0.239 0.0117 0.0909
@@
Media2Face* 0.96 0.308 0.531 0.363 0.0163 0.0459 0.108 0.256 0.0105 0.0899
DiffPoseTalk* 0.68 0.464 0.612 0.267 0.025 0.0379 0.117 0.223 0.0101 0.0776
@@
Table 2: Quantitative evaluation of THUNDER, other models and ablations. (1) Mesh-to-speech input space. All of the mesh-to-speech variants are applicable. THUNDER with mouth2s results in the best lip-sync performance (LVE, L-CCC, L-PCC), which is why we select it as the final THUNDER model. exp2s is applicable, too as its scores best on DTW and has good diversity (S-DIV-U), while face2s scores the worst on diversity (S-DIV-U, T-DIV-U). (2) Trainable audio encoder. THUNDER-T results in improved lip-sync metrics (LVE, L-CCC, L-PCC and DTW), but at the slight expense of diversity of outputs (mainly upper face diversity S-DIV-U and T-DIV-U) compared to THUNDER which does not train the audio encoder. (3) Mesh-To-Speech applied to other methods. The application of mesh-to-speech to our reimplementations of Media2Face* (diffusion with image conditioning), DiffPoseTalk* (diffusion with style feature conditioning) and FlameFormer (deterministic with subject ID conditioning) improves lip-sync metrics at slight expense of expression diversity. Please note that DiffPoseTalk* and Media2Face* (both with and without mesh-to-speech) have a considerably lower diversity than THUNDER. This happens due to the additional conditioning inputs which narrow down the output distribution space. This also manifests itself with better lip-sync performance.
Ablation study. We ablate our design decisions. (1) The input space for the mesh-to-speech loss, (2) the trainability of the audio encoder (Tab. 2) , and (3) the weight of the mesh-to-speech loss (see Sec. 5 in Sup. Mat. and video).
Mesh-to-speech loss with other methods. Tab. 2 shows that mesh-to-speech is applicable to other methods. We re-implement three methods that incorporate additional conditioning mechanisms and train them with and without our mesh-to-speech loss. Our new loss improves performance on the lip-sync metrics. For visual results, refer to Sup. Video. Comparison to additional methods. Tab. 7 in Sup. Mat. we also include a comparison of THUNDER and other (original) implementations recent methods.
Disentanglement effect. To test whether mesh-to-speech helps preserve the effect of good lip-sync in the presence of other editing conditions, we additionally run our reimplementations of Media2Face with conditioning inputs extracted from emotional images. See Sec. 7.2.1 in Sup. Mat.
Other datasets. To verify, whether THUNDER is dataset-agnostic, we also trained and tested the method on TFHP. We report the results in Sec. 7.2.4 in Sup. Mat..
Perceptual study. We conduct a perceptual study in which we evaluate THUNDER and the effect of mesh-to-speech. Specifically, we compare to the deterministic FlameFormer, THUNDER with and without mesh-to-speech, and p-GT. The participants are shown two videos side-by-side and asked to rate animations on a five-point Likert scale (strong/weak preference for A or B or indifferent). We ask the participants to rate three aspects of the animation: lip-sync quality, dynamism, and realism. Fig. 4 reports the results for THUNDER and Fig. 5 for THUNDER-T. For the remaining results of the study and further discussion, see Sec. 13 in Sup. Mat.

### v3

@@ 4.2 Speech-Driven Facial Animation Diffusion
Figure 3: THUNDER architecture. The upper part of the figure (green) depicts the architecture of the diffusion model and the lower part (yellow) illustrates the application of the audio consistency loss. Gray boxes indicate the input to the system. Trainable components are highlighted in orange and frozen ones in blue.
Architecture. The core of THUNDER consists of a denoiser in the form of a transformer decoder [86]. Similar to other contemporary 3D animation methods (e.g. Zhao et al. [98] and Sun et al. [74]), THUNDER processes a noisy sequence of expression parameters $\boldsymbol{\tilde{x}}_{(d)}^{1:T}$, where $d$ represents the diffusion timestep. These parameters serve as the query inputs to the transformer. The keys and values for the transformer are comprised of the input audio features $\mathbf{a}^{1:T}$ and a diffusion step embedding $\mathcal{T}(d)$, concatenated along the temporal dimension. The function $\mathcal{T}$ is defined as a sinusoidal timestep function similar to Ho et al. [36]. We do not incorporate additional animation-controlling conditions (such as style vectors [16, 26, 91], emotion vectors [18, 61] or CLIP embeddings [98]) as this is not the goal of this work. However, the design can easily be augmented with additional input conditions. All inputs are projected into the transformer’s feature dimension, $f=128$, using dedicated learnable affine layers. Positional encoding within the transformer’s self-attention utilizes the ALiBi mechanism [65]. The cross-attention layers use a diagonal binary mask to facilitate the attention mechanism, and no additional positional encoding. The model predicts the reconstructed expression parameters $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. The transformation can be formally described by the following equation:
@@
Here, $d\in\{1,\dots,D\}$ indicates the diffusion timestep, and $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$ is the denoised prediction. Unlike some methods that predict noise, this model directly predicts the fully denoised space, which is used to compute the 3D mesh sequence: $M(\boldsymbol{\beta}=\mathbf{0},\boldsymbol{\widehat{x}}_{(0)}^{1:T})\rightarrow\widehat{\mathbf{V}}_{(0)}^{1:T}$. This approach allows the model to define losses directly on the final 3D mesh space, crucial for leveraging M2S. See Fig. 3 for an overview.
@@
where $\alpha_{d}$ represents the proportion of the original signal’s variance preserved at diffusion timestep $d$, and $\epsilon$ is drawn from a Gaussian distribution with predefined $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$. This approach incrementally corrupts the clean parameters, enabling the model to learn to denoise varying levels of noise. The model is supervised with several loss terms - reconstruction and velocity losses in expression space and 3D vertex space, and also the M2S loss (Eq. (5)). The reconstruction loss is given as:
$\mathcal{L}_{rec}=\ \|\mathbf{V}^{1:T}-\widehat{\mathbf{V}}_{(0)}^{1:T}\|_{2}^{2}+\|\boldsymbol{\psi}^{1:T}-\boldsymbol{\widehat{\psi}}_{(0)}^{1:T}\|_{2}^{2}+\ \|\boldsymbol{\theta}_{jaw}^{1:T}-{\boldsymbol{\widehat{\theta}}_{jaw}^{1:T}}_{(0)}\|_{2}^{2}.$
The velocity loss is also an MSE error but computed on the velocities of the individual components:
$\mathcal{L}_{vel}=$ $\ \|v(\mathbf{V}^{1:T})-v(\widehat{\mathbf{V}}_{(0)}^{1:T})\|_{2}^{2}+\|v(\boldsymbol{\psi}^{1:T})-v(\boldsymbol{\widehat{\psi}}_{(0)}^{1:T})\|_{2}^{2}+$
$\ \|v(\boldsymbol{\theta}_{jaw}^{1:T})-v({\boldsymbol{\widehat{\theta}}_{jaw}^{1:T}}_{(0)})\|_{2}^{2},$
@@
with $w_{m2s}=1$. The model is trained using classifier-free guidance [35]. The audio condition is randomly dropped and replaced with a learnable vector with 20% probability.
Inference. At inference time, we initialize the first step by sampling the Gaussian noise: $\boldsymbol{\widehat{x}}_{(D)}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. We then proceed to denoise for $D$ diffusion steps. In each step, we first employ the denoiser $\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})$ to obtain $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. Then, we compute $\boldsymbol{\tilde{x}}_{(d-1)}$ by adding the expected amount of noise back to obtain $\boldsymbol{\tilde{x}}_{(d-1)}$. Since THUNDER is trained with classifier-free guidance, it is possible to trade-off fidelity of the lip-sync for increased diversity by combining $s_{a}\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})+(1-s_{a})\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\emptyset)$ during the denoising process. However, we find our results are diverse enough without this and so we effectively set $s_{a}=1$.
@@ 5.1 Datasets
The selection of datasets for THUNDER is challenging. 4D scan datasets such as VOCASET [16] provide good GT but are too small to train generative models. In order to have a high-enough quality geometry from which speech can be predicted, We use a collection of public in-the-lab captured video datasets commonly used for lip-reading, namely TCD-TIMIT [34], RAVDESS [53] and GRID [14], totaling 117 subjects, uttering 42140 short sentences, totaling 3535816 frames at 25FPS (over 39 hours). These datasets provide clear imagery without in-the-wild complexity and can be reconstructed by off-the-shelf methods to a satisfactory degree. Following EMOTE [18], we employ the publicly available INFERNO framework [19], which contains a SOTA face reconstruction system [100, 17, 29]. We include an expanded discussion on face tracker selection in Sec. 9 in Sup. Mat. We split each of the datasets by subjects, 70% for training and 15% each for validation and testing. We refer to this data as THUNDERSET. Additionally, we also train and test M2S and THUNDER on a more challenging TFHP [74], which contains sequences in-the-wild and head pose.
@@ 5.2 Mesh-To-Speech
Implementation details. We initialize the conformer and prediction heads with pretrained weights from Choi et al. [12], while the input MLP is trained from scratch. We train the model with the Adam optimizer [49] for 20 epochs and select the checkpoint with the lowest validation loss (usually reached in 5 epochs). The training batch size is 16 and the input sequence is trimmed to 125 frames at 25FPS.
Quantitative comparison. Since the M2S model is the first of its kind, we compare M2S to a recent SV2S model [12] which we finetune on the videos from our training dataset for fair comparison. Note that FLAME does not model the teeth or tongue, which makes the M2S task arguably harder than SV2S. Teeth and tongue, which may be visible on video, are critical for production of certain phonemes, such as labiodentals (/f/, /v/), or alveolar phonemes (/s/, /z/, /t/, /d/, /l/) [50]. Consequently, we do not expect M2S to match the performance of SV2S methods.
Table 1 reports the standard video-to-speech metrics, namely Short-Time Objective Intelligibility [76], Extended STOI [40] to measure the intelligibility of the generated samples, Perceptual Evaluation of Speech Quality (PESQ) [70] (narrow and wide band) to measure the perceptual quality, and Word Error Rate (WER), which we measure with the SOTA ASR network Whisper[67]. As expected, M2S is not as accurate as video-to-speech, but the performance is comparable, suggesting that there is sufficient audio signal in the 3D to be useful for our downstream task. Note that M2S is, however, more accurate than Choi et al.’s original (but not the fine-tuned) model. We experiment with the following input spaces for M2S: selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$ (mouth2s), all vertices $\mathbf{V}$ (face2s), and FLAME expression parameters $\boldsymbol{x}$ (exp2s) and report the results in Tab. 1. Remarkably, the model which takes global expression vectors, performs slightly better than the selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$. Despite that, we choose mouth2s to be our final model, as it results in better lip-sync supervision (discussed in Sec. 5.3 and Tab. 2). Please refer to the Sup. Video for audible comparison and Sec. 7.1 in Sup. Mat. for further discussion on performance.
@@
Mesh-to-speech (M2S) exp2speech $\boldsymbol{x}$ 0.502 0.273 1.257 1.468 0.648
@@
Video-to-speech (SV2S) Choi et al. (finetuned) 0.555 0.348 1.281 1.511 0.437
@@
Table 1: Quantitative comparison of video-to-speech and forms of M2S (mouth2speech, face2speech and exp2speech).
@@ 5.3 Speech-Driven Facial Animation
Implementation details. THUNDER is trained using the Adam optimizer [49] (lr=1e4lr=1e-4) for 260 epochs with batch size 48, sequence length of 70 frames, and $D=1000$ diffusion steps with a linear noise schedule. The transformer has 8 layers with 4 heads and feature dim $f=128$.
Baselines. We compare THUNDER to other speech-driven avatar methods that are also trained on pGT and which output 3DMM parameters (as opposed to vertices). We reimplemented and trained two SOTA methods, namely FlameFormer* and FlameSelfTalk* (adaptations of FaceFormer [26] and SelfTalk [60]), Media2Face* [98] and DiffPoseTalk* [74] (see Sup. Mat. for details). Unless stated otherwise, we keep the pretrained Wav2Vec2 weights frozen. Models that finetune it are denoted with suffix -T (T for trainable). Reimplemented methods are marked with asterisks*.
@@
Lip Vertex Error. Following [69, 74, 91] we report LVE, which calculates the maximum L2 error of all lip vertices for each frame and then computes the average over all frames.
Dynamic Time Warping. We also report the DTW proposed by Thambiraja et al. [84]. First, distances between the mid-points of lower and upper lips are calculated for both the predictions and GT. The two resulting time series are then used to compute the DTW distance.
Lip Correlation Coefficients. LVE is limited to sparse vertices and ignores temporal dynamics, while DTW reduces lip motion to a single distance. To complement these, we treat each mouth-region vertex coordinate as a time series and compute Pearson (synchrony) and Concordance (synchrony and bias/scale agreement) correlation coefficients between pGT and predictions.
Face dynamic deviation. Following CodeTalker [91], we report Face Dynamics Deviation (FDD), which measures the difference between the temporal std.  deviation of the pGT and the predicted vertices, averaged over a set of vertices. We report upper-FDD (FDD-U) and lip-FDD (FDD-L).
Sample diversity. One of the most important aspects of stochastic models is that they generate diverse animations for the same input. Remarkably, previous work has not analyzed this aspect [73, 74, 98]. We compute standard deviations of vertex distances between the pGT and the predictions for the same input at frame $t$. Formally, given a tensor of per-vertex L2 distances $\mathbf{D}\in\mathbb{R}^{N\times S\ \times T\times n_{v}}$, with $N$ being the size of the test dataset and $S=32$ is the number of sampled outputs per input. We compute the standard deviation along the sample dimension $S$. Other dimensions are subsequently averaged. We measure sample diversity for the upper face region (S-Div-U) and the lip region (S-Div-L).
Temporal diversity. We also report the temporal diversity, where the standard deviation is computed over the temporal dimension of $\mathbf{D}\in\mathbb{R}^{N\times S\ \times T\times n_{v}}$ (and averaged across the rest). Again, we calculate the diversity for both lip (T-Div-L) and upper face regions (T-Div-U).
@@ 5.3.1 Quantitative evaluation
Lip-Sync Upper-face Diversity [cm] Lip Diversity [cm] Face Dynamic Dev. [cm]
Experiment Input Model LVE [cm] $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW [cm] $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\downarrow$ T-FDD-U $\downarrow$ T -FDD-L $\downarrow$
(1) THUNDER and mesh-to-speech input space audio only THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
@@
(2) THUNDER and trainable audio enc. audio only THUNDER-T w/o m2s 0.723 0.428 0.623 0.266 0.020 0.0411 0.0656 0.216 0.0118 0.0811
@@
(3) Mesh-to-speech with other diffusion-based methods which use additional input conditions audio, image feature Media2Face* 0.96 0.308 0.531 0.363 0.0163 0.0459 0.108 0.256 0.0105 0.0899
audio, speaker style feat. DiffPoseTalk* 0.68 0.464 0.612 0.267 0.025 0.0379 0.117 0.223 0.0101 0.0776
(4) Mesh-to-speech with deterministic methods and discrete input conditions audio, one-hot speaker ID FlameFormer* 0.809 0.368 0.57 0.291 0.0271 0.0372 0.132 0.239 0.0117 0.0909
@@
audio, one-hot speaker ID FlameSelfTalk* 0.695 0.504 0.698 0.243 0.0275 0.0372 0.108 0.201 0.00983 0.0692
FlameSelfTalk* w/ m2s 0.674 0.518 0.705 0.243 0.029 0.039 0.108 0.193 0.00965 0.0674
@@
Table 2: Quantitative evaluation on THUNDERSET (1) Mesh-to-speech input space. All of the M2S variants are applicable. THUNDER with mouth2s results in the best lip-sync performance (LVE, L-CCC, L-PCC), which is why we select it as the final THUNDER model. exp2s is applicable, too as it scores best on DTW and has good diversity (S-DIV-U), while face2s scores the worst on diversity (S-DIV-U, T-DIV-U). (2) Trainable audio encoder. THUNDER-T results in improved lip-sync metrics (LVE, L-CCC, L-PCC and DTW), but at slight expense of diversity of outputs (mainly upper face diversity S-DIV-U and T-DIV-U) compared to THUNDER which does not train the audio encoder. (3) Mesh-To-Speech with diffusion-based methods. Training Media2Face* (with image conditioning) and DiffPoseTalk* (style conditioning) follows the same trend - increase in lip-sync quality and reduction in upper-face diversity. (4) Mesh-To-Speech with deterministic methods. We train FlameFormer* and SelfTalk* with M2S and find that it improves the lip-sync related metrics.
Tab. 2 reports the test-set metrics computed for models trained on THUNDERSET and Tab. 3 does so for models trained on TFHP. We analyze the following:
Does M2S improve THUNDER? Yes, models with M2S improve all lip-sync metrics at slight expense of diversity (see Tab. 2, experiments 1 and 2).
Does the input space for M2S matter? While all models are improve lip-sync metrics, mouth2speech has the strongest effect on lip-sync metrics, likely thanks to its localized effect on the mouth region (Tab. 2, experiment 1).
Should the audio encoder be finetuned? Trainable encoder (what most methods use) results in much lower diversity, presumably due to overfitting to the training voices. Frozen encoder results in inferior lip-sync but higher diversity. M2S improves lip-sync in both cases. M2S applied with a frozen encoder leads to a dramatically improved lip-sync while retaining some of the upper-face diversity (Tab. 2, exp. 1,2).
Does M2S improve other methods? Yes, we apply M2S to train the deterministic FlameFormer, SelfTalk and stochastic diffusion-based Media2Face and DiffPoseTalk and show improvement on the lip-sync metrics (Tab. 2, exp. 3,4) .
Does THUNDER generalize to other datasets? Yes. We train and test M2S and THUNDER on the challenging TFHP (the original DiffPoseTalk dataset). THUNDER outperforms original DiffPoseTalk (see Tab. 3 and Sec. 7.2.1 in Sup. Mat.)
Can THUNDER generate head pose motion? Yes. We train a THUNDER model with head pose on TFHP and show improvement over DiffPoseTalk (see Tab. 3 and Sec. 7.2.1).
Does M2S help with editing methods? Yes. We feed images of different emotions to our M2F. M2F w/ M2S improves lip-sync (see Sec. 7.2.2 in Sup. Mat.)
Lip-Sync Upper-face diversity [mm] Lip diversity [mm] Face Dynamic Dev. [mm] Head Pose
Name Backbone LVE [cm] $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW [cm] $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\downarrow$ FDD-U $\downarrow$ FDD-L $\downarrow$ BA $\times 10^{-3}$ $\uparrow$ DIV $\uparrow$
THUNDER-F w/o m2s Wav2Vec2 frozen 1.2 0.285 0.388 2.26 0.258 0.349 2.6 3.58 0.0649 0.76 6.02 3.07
THUNDER-F Wav2Vec2 frozen 1.07 0.298 0.435 1.95 0.203 0.321 1.5 2.98 0.0707 0.95 6.13 3.78
DiffPoseTalk HuBERT trainable 1.01 0.414 0.541 1.85 0.247 0.352 1.58 3.02 0.0641 0.618 5.54 2.16
THUNDER-T w/o m2s Wav2Vec2 trainable 1.12 0.358 0.488 1.86 0.241 0.351 1.87 3.21 0.0597 0.661 6.25 3.49
THUNDER-T Wav2Vec2 trainable 0.987 0.422 0.55 1.81 0.244 0.341 1.35 2.89 0.0578 0.55 5.77 4.83
Table 3: Quantitative evaluation on TFHP. We compare THUNDER, and THUNDER w/o M2S and the original DiffPoseTalk release (all trained on TFHP). Consistently to our other experiments, THUNDER results in superior lip-sync performance over THUNDER w/o M2S and DiffPoseTalk, at a slight trade-off of diversity metrics. Notably, when trained with head pose, all THUNDER variants outperforms DiffPoseTalk on head pose metrics - beat alignment (BA) and head pose diversity (DIV). Refer to Sup. Video and PDF for qualitative results.